OpenStudy (anonymous):

solve for x. show the steps leading to the solution 4(x^2+6x-3)=24-8x+2(2x^2-7)

OpenStudy (mimi_x3):

4x^2 + 24x - 12 = 24 - 8x + 4x^2 - 14 32x = 24+12 - 14 32x = 22 x = 22/32 = 11/16